### v1

@@ Abstract
Mobile app review analysis presents unique challenges due to the low quality, subjective bias, and noisy content of user-generated documents. Extracting features from these reviews is essential for tasks such as feature prioritization and sentiment analysis, but it remains a challenging task. Meanwhile, encoder-only models based on the Transformer architecture have shown promising results for classification and information extraction tasks for multiple software engineering processes. This study explores the hypothesis that encoder-only large language models can enhance feature extraction from mobile app reviews. By leveraging crowdsourced annotations from an industrial context, we redefine feature extraction as a supervised token classification task. Our approach includes extending the pre-training of these models with a large corpus of user reviews to improve contextual understanding and employing instance selection techniques to optimize model fine-tuning. Empirical evaluations demonstrate that this method improves the precision and recall of extracted features and enhances performance efficiency. Key contributions include a novel approach to feature extraction, annotated datasets, extended pre-trained models, and an instance selection mechanism for cost-effective fine-tuning. This research provides practical methods and empirical evidence in applying large language models to natural language processing tasks within mobile app reviews, offering improved performance in feature extraction.
@@ 1 Introduction
In the context of Requirements Engineering (RE), effective adoption and proper use of LLMs have been elicited as key challenges for future work in the field (Frattini et al., 2024; Ronanki et al., 2023; Fantechi et al., 2023). As a text-based document-driven community, both generated by developers (e.g., textual requirements, user stories, test cases) and by users (e.g., bug reports, software issues, app reviews), information extraction from these documents is key to support requirements elicitation (Ronanki et al., 2023), defect repair (Ferrari et al., 2018), prioritization (Malgaonkar et al., 2022), information extraction (Sleimi et al., 2021), feedback analysis (Dalpiaz and Parente, 2019), and release planning (McZara et al., 2015; Sharma and Kumar, 2019). Extracting descriptors, metadata, entities and categories from these documents allows practitioners to categorize large amounts of data and build analytical processes, which is especially useful for continuously analysing large amounts of user-generated documents (Li et al., 2018; van Vliet et al., 2020). A particular example are opinion mining methods for mobile app reviews (Dabrowski et al., 2022), for which three major tasks can be identified from the literature: (1) review classification (e.g., bug report, feature request (Maalej and Nabil, 2015)); (2) sentiment analysis (e.g., positive, negative (Zhang et al., 2014)); and (3) app feature extraction (e.g., send message, make video call (Johann et al., 2017)). Concerning the latter, the automatic extraction of mobile app features (i.e., functions or characteristics of a mobile app from the user perspective) supports multiple feature-oriented decision-making tasks, including feature prioritization (Scalabrino et al., 2019) and feature-oriented sentiment analysis (Guzman and Maalej, 2014).
While some feature extraction methods have been proposed by leveraging syntactic-based pattern matching techniques (Johann et al., 2017; Guzman and Maalej, 2014; Dragoni et al., 2019), several challenges remain (Dabrowski et al., 2022). For starters, agreement towards what constitutes a feature is typically low, as it is considered a particularly cognitively subjective task (Dabrowski et al., 2023). In addition, app reviews are low-quality user-generated documents, subjectively biased, grammatically incorrect, relatively short, filled with noisy content and even potentially generated by bots (Araujo et al., 2022). Hence, traditional and machine learning (ML) methods for feature extraction typically struggle especially in the context of app reviews, reporting an overall low recall and hence missing multiple feature mentions from a large corpus of reviews (Dabrowski et al., 2023).
In this context, this study aims to validate the following hypothesis:
@@
Encoder-only LLMs utilize only the encoder component of the Transformer architecture without a decoder (Minaee et al., 2024), making them suitable for tasks like classification, named-entity recognition (NER) and information extraction. To this end, our approach leverages crowdsourced user annotations from an industrial context to redefine the feature extraction task as a supervised token classification (e.g. NER) task. We compare the performance of multiple encoder-only LLMs with a classification layer on top to extract subsets of tokens referring to particular features for a given mobile app.
Using this approach as a baseline, we complement our research with two additional hypotheses to explore the potential of improving the functional suitability and performance efficiency of our LLM-based approach:
@@
    Mobile app user review context can be better reflected in general-purpose LLMs by extending the pre-training with a large corpus of user reviews, improving the functional suitability of feature extraction.
@@
    Instance selection of crowdsourced user reviews can optimize model fine-tuning, improving the performance efficiency of feature extraction while maintaining - or even improving - functional suitability.
@@
    A document-based instance selection mechanism for token classification tasks to support cost-effectiveness assessment of fine-tuning LLMs for NER.
[…]
    An empirical analysis of the suitability and the performance variations across the combination of (1) different types of encoder-only LLMs, (2) extended pre-training settings, and (3) instance selection data partitions.
@@ 2.1 Feature extraction
Features have become a core descriptor for app review mining activities (Dabrowski et al., 2022). Furthermore, features are also used to support user-oriented services, including mobile app recommendation (Palomba et al., 2015), search-based algorithms (Chen et al., 2016) and personalization (Laranjo et al., 2021). Figure 1 illustrates a practical example of features being used as indexation descriptors to support alternative software recommendations in AlternativeTo [^2] Source: https://alternativeto.net/software/telegram/about/. Features are used to cluster and identify similar mobile apps and provide recommendations based on similar features. Additionally, rare features are also highlighted to illustrate the app’s singularity. However, these features are extracted based on manual user votes, limiting their generalization, and causing biased, unbalanced representativity and incomplete feature data.
Alternatively to manual annotation, feature extraction refers to the automatic identification of features mentioned within a textual document, including app descriptions and user reviews (Dabrowski et al., 2022). While formal methods are mostly based on syntactic-based pattern-matching and topic modelling (see Section 7), several challenges remain. From a formalization perspective, scoping and stating what constitutes a feature is highly subjective, leading to discrepancies affecting methodological developments and even different application scenarios between industrial and research environments (Johann et al., 2017). From a methodological perspective, due to the nature and authorship of reviews, these methods tend to struggle when processing noisy, grammatically inaccurate and complex documents (Shah et al., 2019). From an evaluation point of view, this leads to low recall values, producing high false negatives and therefore limiting the added value of an automatic approach by missing multiple feature candidates (Dabrowski et al., 2023).
@@ 2.2 Token classification and NER
Token classification is a fundamental task in the NLP field where the goal is to assign a particular label (i.e., class) to each token within a text sequence (Naveed et al., 2024). In this context, given a review $r$, let $T(r)$ denote the sequence of tokens in $r$. Formally, if $r$ consists of $n$ tokens, then $T(r)=[t_{1},t_{2},\ldots,t_{n}]$, where $t_{i}$ represents the $i$-th token in the sequence.
Let $C$ be the set of possible classes for each token. The objective of token classification is to find a function $\phi:T(r)\rightarrow C$ that maps each token $t_{i}$ to a class $c_{i}\in C$. This can be expressed as:
$\phi(t_{i})=c_{i},\quad\forall t_{i}\in T(r)$
The annotated sequence for a review $r$ is thus $[(t_{1},c_{1}),(t_{2},c_{2}),\ldots,(t_{n},c_{n})]$.
NER is a specialized form of token classification where the objective is to identify and classify specific entities within a text into predefined categories such as names of people, organizations, locations, dates, and other domain-specific entities (Hou et al., 2024). Each token in a sequence is assigned a label that indicates whether it is part of an entity and its role within that entity, such as the beginning ($B-$), inside ($I-$), or outside ($O$) of a named entity.
@@
To address the limitations of syntactic-based approaches, and based on the validation of H1, we propose using encoder-only LLMs fine-tuned for function $\phi$ through a supervised learning approach. This method leverages crowdsourced annotations generated by real users in an industrial context (i.e., AlternativeTo).
@@ 2.3 LLMs and extended pre-training
To enhance the contextual understanding of these models in the mobile app review domain, and based on the validation of H2, we propose extending the pre-training of encoder-only LLMs with a large dataset of reviews from various popular mobile app categories. This aims to improve the functional suitability of the feature extraction $\phi$ method.
@@ 2.4 Instance selection
Instance selection involves identifying and retaining the most relevant documents (e.g., mobile app reviews) from a corpus while filtering out non-relevant or redundant instances (Cunha et al., 2023). This technique aims to enhance the efficiency and effectiveness of machine- and deep -learning models. Efficiency can be improved by reducing large datasets filtering non-relevant documents (i.e., documents not affecting the task performance in terms of functional suitability), especially when training on the entire corpus may be computationally prohibitive (Wilson and Martinez, 2000). Additionally, effectiveness can be improved by removing document instances that might be noisy, mislabeled, or highly redundant (Carbonera, 2017). There are multiple instance selection strategies and methods according to the goal for filtering document instances. This includes removal of mislabeled (Wilson, 1972) or noisy (Wilson and Martinez, 2000) documents using gold-standards, selection of highly representative documents through density measures (Carbonera and Abel, 2015), and clustering-based approaches (Moran et al., 2022).
@@
To optimize resource consumption for fine-tuning methods, and based on the validation of H3, we explore the impact of instance selection techniques based on redundancy reduction and increased representativeness of domain-specific knowledge. This approach aims to reduce the required dataset for fine-tuning while maintaining, or even improving, functional quality.
@@ 3 Study design
3.1 Objective and research questions
@@ 3.2 Research method
Figure 2 illustrates a general overview of the research method conducted in this study. Following a Design Science (DS) methodology for software engineering research (Engström et al., 2020), we refined the main goal of this research into three scientific sub-objectives focusing on the validation of H1 $\rightarrow$ H3. We aligned each Design Science iteration with a particular sub-objective, leading to three iterations covering (i) objective refinement, (ii) design and development of the solution, and (iii) empirical evaluation and verification of results. Concerning evaluation, we use the ISO/IEC 25010 software product quality model (International Organization for Standardization, 2023) to focus on two quality characteristics for the evaluation: functional suitability, for which we focus on functional correctness with respect to ground truth data and human evaluation; and performance efficiency, for which we focus on time behaviour. Extended details of the evaluation design are depicted in Section 5.
Figure 2: Research method based on DS for software engineering research
@@ 4.1.1 Data collection and annotation
  • Features dataset. Using the set of reviews, we implemented web scraping mechanisms to access archived cached versions from AlternativeTo using the Wayback Machine Internet Archive project [^5] http://web.archive.org/ for automatically extracting feature annotations by users for each app in the original dataset. This process resulted in 198 distinct features for 468 mobile apps.
@@ 4.1.2 Data pre-processing and feature transfer
For review and feature pre-processing 2, we apply a common natural language pre-processing stage using Stanza’s neural pipeline [^6] https://stanfordnlp.github.io/stanza/neural_pipeline.html for syntactic annotation and entity extraction from user reviews and features. The pipeline was composed of the following steps: (i) tokenization, (ii) multi-word token expansion, (iii) PoS tagging, (iv) morphological feature extraction, and (v) lemmatization. The output of this pipeline is formatted using the CoNLL-U format [^7] https://universaldependencies.org/format.html. After the pre-processing step, the corpus is ready for the annotation process, which we depict in detail in Algorithm 1. Each token $t\in T(r),\forall r\in R$ is initialized with the default label $O$. The algorithm looks for all matches between feature tokens and review tokens. If a match is found for a particular review $r$ and feature $f$ for a given mobile app $app$, which means that users from AlternativeTo voted $f$ as a feature from $app$, then each token resulted from the intersection $T(r)\cap T(f)$ is annotated with B-feature or I-feature according to the position of the token within the original feature.
@@ 4.1.3 Model fine-tuning
Table 1: Model features and fine-tuning parameters.
@@ 5.1 Design
5.1.1 Functional suitability - Ground truth
For functional correctness, we use collected reviews from Google Play and features from AlternativeTo as ground truth by assessing the correctness of $\phi(t_{i})$ predictions $\forall t_{i}\in T(r)$ and $\forall r\in R$, where $R$ is the corpus of reviews used for evaluation (see Section 5.2). Based on model predictions, we focus on the following metrics for measuring functional suitability. First, we focus on precision:
$p=TP/(TP+FP)$
@@ 5.1.1 Functional suitability - Ground truth
Next, we focus on recall:
$r=TP/(TP+FN)$
where $FN$ refers to tokens with a predicted class $c_{i}=O$ not matching the ground truth class. Recall measures how many feature tokens were missed by the feature extraction method.
Then, we focus on the f-measure:
$f_{1}=2\cdot\frac{p\cdot r}{p+r}$
which measures the harmonic mean between precision and recall.
@@
5.1.2 Functional suitability - Human evaluation
@@
The human evaluation process consisted of the following steps:
@@
    Figure 6: Questionnaire for automatic feature extraction ($Q_{A}$).
    Figure 7: Questionnaire for manual feature extraction ($Q_{M}$).
@@
      F-measure weighting factor $\beta$ ($Q_{A}$ and $Q_{M}$). We randomly selected a subset of 100 reviews from the evaluation set. For this subset, we created one $Q_{A}$ questionnaire with 100 reviews and four $Q_{M}$ questionnaires with 25 reviews each. We measured the time required for each annotator to complete the questionnaire, and we report average values among all valid annotators to measure $A_{t}$ and $A_{T}$.

### v2

@@ Abstract
Mobile app review analysis presents unique challenges due to the low quality, subjective bias, and noisy content of user-generated documents. Extracting features from these reviews is essential for tasks such as feature prioritization and sentiment analysis, but it remains a challenging task. Meanwhile, encoder-only models based on the Transformer architecture have shown promising results for classification and information extraction tasks for multiple software engineering processes. This study explores the hypothesis that encoder-only large language models can enhance feature extraction from mobile app reviews. By leveraging crowdsourced annotations from an industrial context, we redefine feature extraction as a supervised token classification task. Our approach includes extending the pre-training of these models with a large corpus of user reviews to improve contextual understanding and employing instance selection techniques to optimize model fine-tuning. Empirical evaluations demonstrate that these methods improve the precision and recall of extracted features and enhance performance efficiency. Key contributions include a novel approach to feature extraction, annotated datasets, extended pre-trained models, and an instance selection mechanism for cost-effective fine-tuning. This research provides practical methods and empirical evidence in applying large language models to natural language processing tasks within mobile app reviews, offering improved performance in feature extraction.
@@ 1 Introduction
In the context of Requirements Engineering (RE), effective adoption and proper use of LLMs have been elicited as key challenges for future work in the field (Frattini et al., 2024; Ronanki et al., 2023; Fantechi et al., 2023). As a text-based document-driven community, both generated by developers (e.g., textual requirements, user stories, test cases) and by users (e.g., bug reports, software issues, app reviews), information extraction from these documents is key to support requirements elicitation (Ronanki et al., 2023), defect repair (Ferrari et al., 2018), prioritization (Malgaonkar et al., 2022), information extraction (Sleimi et al., 2021), feedback analysis (Dalpiaz and Parente, 2019), and release planning (McZara et al., 2015; Sharma and Kumar, 2019). Extracting descriptors, metadata, entities and categories from these documents allows practitioners to categorize large amounts of data and build analytical processes, which is especially useful for continuously analysing large amounts of user-generated documents (Li et al., 2018; van Vliet et al., 2020). A particular example are opinion mining methods for mobile app reviews (Dabrowski et al., 2022), for which three major tasks can be identified from the literature: (1) review classification (e.g., bug report, feature request (Maalej and Nabil, 2015)); (2) sentiment analysis (e.g., positive, negative (Zhang et al., 2014)); and (3) app feature extraction (e.g., send message, make video call (Johann et al., 2017)). Concerning the latter, the automatic extraction of mobile app features (i.e., functions or quality characteristics of a mobile app from the user perspective) supports multiple feature-oriented decision-making tasks, including feature prioritization (Scalabrino et al., 2019) and feature-oriented sentiment analysis (Guzman and Maalej, 2014).
While some feature extraction methods have been proposed by leveraging syntactic-based pattern matching techniques (Johann et al., 2017; Guzman and Maalej, 2014; Dragoni et al., 2019), several challenges remain (Dabrowski et al., 2022). For starters, agreement towards what constitutes a feature is typically low, as it is considered a particularly cognitively subjective task (Dabrowski et al., 2023). In addition, app reviews are low-quality user-generated documents, subjectively biased, grammatically incorrect, relatively short, filled with noisy content and even potentially generated by bots (Araujo et al., 2022). Hence, syntactic-based and machine learning (ML) methods for feature extraction typically struggle especially in the context of app reviews, reporting an overall low recall and hence missing multiple feature mentions from a large corpus of reviews (Dabrowski et al., 2023).
To address these challenges, encoder-only LLMs, which rely solely on the encoder component of the Transformer architecture (Minaee et al., 2024), have emerged as a promising solution. These models are particularly well-suited for tasks such as classification, named-entity recognition (NER), and information extraction. In this context, this study aims to validate the following hypothesis:
@@
To this end, our approach leverages crowdsourced user annotations from an industrial context to redefine the feature extraction task as a supervised token classification (e.g. NER) task. We compare the performance of multiple encoder-only LLMs with a classification layer on top to extract subsets of tokens referring to particular features for a given mobile app.
Using this approach as a baseline, we complement our research with two additional hypotheses to explore the potential of improving the functional correctness and time behaviour of our LLM-based approach:
@@
    Extending the pre-training of general-purpose LLMs (i.e., additional pre-training on domain-specific data) with a large corpus of user reviews can improve the representation of the mobile app user review context, improving the functional correctness of feature extraction.
@@
    Instance selection of crowdsourced user reviews (i.e., filtering and selecting relevant instances for fine-tuning) can optimize model fine-tuning, improving the time behaviour of feature extraction while maintaining - or even improving - functional correctness.
@@
    A document-based instance selection method for token classification tasks to support cost-effectiveness assessment of fine-tuning LLMs for NER.
[…]
@@
    An empirical analysis of the correctness and the execution time variations across the combination of (1) different types of encoder-only LLMs, (2) extended pre-training settings, and (3) instance selection data partitions.
@@ 2.1 Feature extraction
Features have become a core descriptor for app review mining activities (Dabrowski et al., 2022). Furthermore, features are also used to support user-oriented services, including mobile app recommendation (Palomba et al., 2015), search-based algorithms (Chen et al., 2016) and personalization (Laranjo et al., 2021). Figure 1 illustrates a practical example of features being used as categorization tags to support alternative software recommendations in AlternativeTo [^2] Source: https://alternativeto.net/software/telegram/about/. Features are used to cluster and identify similar mobile apps and provide recommendations based on similar features. Additionally, rare features are also highlighted to illustrate the app’s singularity. However, these features are extracted based on manual user votes, limiting their generalization, and causing biased, unbalanced representativity and incomplete feature data.
Feature extraction refers to the automatic identification of features mentioned within a textual document, including app descriptions and user reviews (Dabrowski et al., 2022). While formal methods are mostly based on syntactic-based pattern-matching and topic modelling (see Section 7), several challenges remain. From a formalization perspective, scoping and stating what constitutes a feature is highly subjective, leading to discrepancies affecting methodological developments and even different application scenarios between industrial and research environments (Johann et al., 2017). From a methodological perspective, due to the nature and authorship of reviews, these methods tend to struggle when processing noisy, grammatically inaccurate and complex documents (Shah et al., 2019). From an evaluation point of view, this leads to low recall values, producing high false negatives and therefore limiting the added value of an automatic approach by missing multiple feature candidates (Dabrowski et al., 2023).
@@ 2.2 Token classification and NER
Token classification is a fundamental task in the NLP field where the goal is to assign a particular label (i.e., class) to each token within a text sequence (Naveed et al., 2024). Given a review $r$, let $T(r)$ denote the sequence of tokens in $r$. Formally, if $r$ consists of $n$ tokens, then $T(r)=[t_{1},t_{2},\ldots,t_{n}]$, where $t_{i}$ represents the $i$-th token in the sequence. Hence, token classification involves assigning one of the possible classes from a predefined set $C$ to each token $t_{i}$ in the sequence. As a result, the annotated sequence consists of pairs of tokens and their assigned classes, such as $[(t_{1},c_{1}),(t_{2},c_{2}),\ldots,(t_{n},c_{n})]$, where each $c_{i}$ represents the label assigned to the corresponding token $t_{i}$.
NER is a specialized form of token classification where the goal is to identify and classify specific entities within a text into predefined categories such as names of people, organizations, locations, dates, and other domain-specific entities (Hou et al., 2024). This is achieved using the IOB tagging scheme, where each token in a sequence is assigned a label that indicates whether it is at the beginning ($B-$), inside ($I-$), or outside ($O$) of a named entity.
@@ 2.4 Instance selection
Instance selection involves identifying and retaining the most relevant documents (e.g., mobile app reviews) from a corpus while filtering out non-relevant or redundant instances (Cunha et al., 2023). This technique aims to enhance the efficiency and effectiveness of machine- and deep -learning models. Efficiency can be improved by reducing large datasets filtering non-relevant documents (i.e., documents not affecting the task performance in terms of functional correctness), especially when training on the entire corpus may be computationally prohibitive (Wilson and Martinez, 2000). Additionally, effectiveness can be improved by removing document instances that might be noisy, mislabeled, or highly redundant (Carbonera, 2017). There are multiple instance selection strategies and methods according to the goal for filtering document instances. This includes removal of mislabeled (Wilson, 1972) or noisy (Wilson and Martinez, 2000) documents using gold-standards, selection of highly representative documents through density measures (Carbonera and Abel, 2015), and clustering-based approaches (Moran et al., 2022).
@@ 3 Study design
3.1 Goal and research questions
@@ 3.2 Research method
Figure 2 illustrates a general overview of the research method conducted in this study. Following a Design Science (DS) methodology for software engineering research (Engström et al., 2020), we refined the main goal of this research into three scientific sub-objectives focusing on the validation of H1 $\rightarrow$ H3. We aligned each Design Science iteration with a particular sub-objective, leading to three iterations covering (i) objective refinement, (ii) design and development of the solution, and (iii) empirical evaluation and verification of results. Concerning evaluation, we use the ISO/IEC 25010 software product quality model (International Organization for Standardization, 2023) to focus on two quality characteristics for the evaluation: functional correctness with respect to ground truth data and human evaluation; and time behaviour of fine-tuning processes. Extended details of the evaluation design are depicted in Section 5.
Figure 2: Research method
@@ 4.1.1 Data collection and annotation
  • Features dataset. Using the set of reviews, we implemented web scraping mechanisms to access archived cached versions from AlternativeTo using the Wayback Machine Internet Archive project [^5] http://web.archive.org/ to automatically extract feature annotations by users. Using the BeautifulSoup library [^6] https://pypi.org/project/beautifulsoup4/, we dynamically inspect the AlternativeTo profile page of each app of the reviews dataset, and we locate and parse the ‘Features’ section (as shown in Figure 1) to extract a flat list of features for each app (e.g., for Telegram, we collect features such as ‘Encrypted Chat’, ‘Large File Transfer’ and ‘Animated stickers’). This resulted in 198 distinct features for 468 apps.
@@ 4.1.2 Data pre-processing and feature transfer
Algorithm 1 depicts in detail the steps for transferring the features collected from AlternativeTo into annotations within the dataset of reviews. For review and feature pre-processing 2, we apply a common natural language pre-processing stage using Stanza’s neural pipeline [^7] https://stanfordnlp.github.io/stanza/neural_pipeline.html for syntactic annotation and entity extraction from both the reviews and features datasets (lines 1-3). The pipeline was composed of the following steps: (i) tokenization, (ii) multi-word token expansion, (iii) PoS tagging, (iv) morphological feature extraction, and (v) lemmatization. The output of this pipeline is formatted using the CoNLL-U format [^8] https://universaldependencies.org/format.html. After the pre-processing step, the corpus is ready for the annotation process. Each token $t\in T(r),\forall r\in R$ is initialized with the default label $O$ (lines 5-7). The algorithm looks for all matches between feature tokens and review tokens (lines 9-21). If a match is found for a particular review $r$ and feature $f$ for a given mobile app $app$ (lines 11-12), which means that users from AlternativeTo voted $f$ as a feature from $app$, then each token resulted from the intersection $T(r)\cap T(f)$ is annotated with B-feature or I-feature according to the position of the token within the original feature (lines 13-18).
@@ 4.1.3 Model fine-tuning
Table 1: Model features and fine-tuning parameters
@@ 5.1 Design
5.1.1 Functional correctness - Ground truth
For functional correctness, we use collected reviews from Google Play and features from AlternativeTo as ground truth by assessing the correctness of $\phi(t_{i})$ predictions $\forall t_{i}\in T(r)$ and $\forall r\in R$, where $R$ is the corpus of reviews used for evaluation (see Section 5.2). Based on model predictions, we focus on the following metrics for measuring functional correctness: precision, recall and f-measure.
@@ 5.1.1 Functional correctness - Ground truth
  • Recall ($r$) measures the proportion of correctly predicted feature-related tokens compared to all feature-related tokens in the ground truth, highlighting how many features were successfully captured by the model.
  • F-measure ($f$) provides a balanced evaluation by calculating the harmonic mean of precision and recall, ensuring that both false positives and missed features are taken into account.
@@
5.1.2 Functional correctness - Human evaluation
@@
    Figure 6: Questionnaire for automatic feature extraction ($Q_{A}$)
    Figure 7: Questionnaire for manual feature extraction ($Q_{M}$)
@@
      F-measure weighting factor $\beta$ ($Q_{A}$ and $Q_{M}$). We randomly selected a subset of 100 reviews from the evaluation set. For this subset, we created one $Q_{A}$ questionnaire with 100 reviews and four $Q_{M}$ questionnaires with 25 reviews each. We measured the time required for each annotator to complete the questionnaire, and we report average values among all valid annotators to measure $A_{t}$ and $A_{T}$, which were then used to compute $\beta=A_{T}/A_{t}$.